# CaMML: Context-Aware Multimodal Learner for Large Models

### v2

@@ Abstract
In this work, we introduce Context-Aware MultiModal Learner (CaMML), for tuning large multimodal models (LMMs). CaMML, a lightweight module, is crafted to seamlessly integrate multimodal contextual samples into large models, thereby empowering the model to derive knowledge from analogous, domain-specific, up-to-date information and make grounded inferences. Importantly, CaMML is highly scalable and can efficiently handle lengthy multimodal context examples owing to its hierarchical design. Based on CaMML, we have developed two multimodal models, CaMML-7B and CaMML-13B, that have shown exceptional performance across an array of benchmark datasets for multimodal tasks. Remarkably, CaMML-13B achieves the state-of-the-art performance on over ten widely recognized multimodal benchmark datasets, surpassing LLaVA-1.5 (13B) with a noticeable margin, without integration of any external resources. Moreover, we have conducted extensive ablative studies to inspect the inner workings of CaMML and performed qualitative analyses to showcase its effectiveness in handling real-world challenging cases.
@@ 1 Introduction
Recently, large multimodal models (LMMs) [32, 61, 54, 31, 66, 53, 3] have demonstrated remarkable performance in a variety of tasks, including but not limited to visual question answering, image captioning, visual grounding, visual-language reasoning, optical character recognition, and visual entailment. Notably, in certain benchmark assessments, these multimodal foundation models have even exceeded human-level performance [63, 32].
Despite the impressive performance, their ability to make inferences is constrained by the knowledge encoded in the model parameters. The inflexible design of these models makes it challenging for them to generalize from contextual examples. For instance, LLaVA-1.5 falls short in processing multiple images and attributes it to the lack of corresponding instruction-tuning training data [31]. Nevertheless, learning and making inferences through contextual examples are fundamental elements of our cognitive processes. Human beings frequently tackle intricate problems by relying on past experiences and identifying analogous situations. Taking inspiration from the cognitive process, we hypothesize that empowering large multimodal models with the capability to perceive and derive insights from analogous contextual examples can significantly streamline the inference process and lead to more precise predictions. Nonetheless, the means to replicate the human cognitive processes in LMMs remain unclear.
As such, our goal is to empower multimodal foundational models to harness context-aware learning, thereby enhancing their ability to comprehend and adapt to previously unseen examples. Identifying relevant context examples is relatively straightforward; this process can be facilitated using multimodal embedding models like ImageBind [15] or CLIP [43]. This approach simulates the act of recalling similar situations from past experiences. Yet, effectively and efficiently integrating the identified similar context samples into large models poses challenges, particularly given the potential variability in the number of context samples and interleaved modalities, resulting in lengthy and heterogeneous context input.
@@ 2 Related Work
The success of LLMs has sparked a burgeoning interest in scaling up multimodal models. One prevalent strategy involves the incorporation of vision encoders, such as ViT [13] and CLIP, into existing LLMs (e.g., LLaMA [52], Vicuna [65]). For example, LLaMA-adapter [61] inserts learnable adaption prompts into LLaMA and make it to follow multimodal instructions for multimodal reasoning. BLIP2 [28] bridges different modalities via Q-Former and employs a two-stage training method to bootstrap both representation learning and LLM (e.g., OPT [62], Flan-T5 [11]) generative learning capabilties. LLaVA [32, 31] projects encoded image features to text token space using linear layers and shows state-of-the-art performance on a variety of multimodal tasks. Another noteworthy series of strategy is to unify multimodal input data and pretrain the model from scratch, exemplified by OFA [54], Perceiver [19, 20], Uni-Perceiver [67], Unival [47], and Unified-IO [35]. These models map images, text, and other modalities into the same IO space and use a causal language model objective for model training.
@@
Extensive research has been conducted on learning from a limited number of multimodal examples. Among them, Flamingo [3] introduces gated xatten-dense layers to establish cross-modal interactions between visual input and text input for few-shot learning. Frozen [53] trains a vision encoder to produce a sequence of image embeddings and input it to frozen language models for multimodal tasks. An alternative approach for tackling multimodal few-shot learning involves leveraging retrieval augmented generations (RAG) [26, 7, 44, 23]. This technique enables the models to access external knowledge repositories, databases, or structured data sources, allowing them to access the up-to-date as well as domain-specific expertise when crafting responses. Recently, there has been a surging interest for harnessing RAG within the realm of multimodal models [9, 55, 56, 33]. Amalgamating RAG with multimodal models helps provide contextually-rich responses and opens up exciting possibilities across domains such as image captioning [55], image generation [56], etc. Among them, Chen et al. [9] propose MuRAG, a model of size 527M, for multimodal-QA. MuRAG concatenates the visual embeddings (extracted by ViT) and text word embeddings of retrieved multimodal image-text pairs and requires the language model to handle a very lengthy input sequences. Re-ViLM [55] is tailored for image captioning and it follows the Flamingo architecture and only uses the retrieved text for augmentation. RA-CM3 [56] leverages retrieval to augment the CM3 [2] pipeline which can do both text-to-image and image-to-text generations.
In contrast to previous approaches, our primary objective is to develop a lightweight module capable of processing multimodal context information efficiently and effectively, with superior generalization capabilities. This novel approach enables LLM to proficiently reason from multimodal in-context examples, leading to more accurate and precise inferences. As discussed in [31], LLaVA-1.5 faces limitations when confronted with scenarios involving multiple images and lengthy contexts. CaMML can effectively address these challenges, a fact further corroborated by its superior performances across multiple benchmarks.
@@ 3.1.1 Datastore and Context Retriever
The datastore is created from either the training set or external resources. We adopt an embedding encoder, ImageBind [15], to extract dense vector representations for every multimodal sample contained in the datastore. Following this, we build an index using Faiss [22], a highly efficient similarity search library, to enable rapid and efficient search operations. During both the training and inference phases, our approach involves identifying the top $N$ most closely related multimodal samples, denoted as $C_{1},C_{2},\ldots,C_{N}$, from the datastore using Faiss, based on a given multimodal query $q$ [^1] Specifically, we compute the similarity between query text and context image, and top-k images along with their corresponding texts are recalled from the datastore.. It is worth noting that the retriever remains frozen throughout all stages of the process.
@@ 3.1.2 Multimodal CaMML Perceiver
To seamlessly integrate interleaved multimodal samples, we introduce a novel module denoted as CaMML Perceiver, which revolves around two key design principles: (1) The module should accommodate a dynamic number of context samples without imposing a significant computational burden. It is worth noting that directly concatenating the images tokens and text tokens will result in $L=\sum^{N}_{i=1}(T_{i}^{\text{IMG}}+T_{i}^{\text{TXT}})$ tokens, where $T_{i}^{\text{IMG}}$ is the number of image tokens and $T_{i}^{\text{TXT}}$ is the number of text tokens. The formulation results in a linear increase of $L$ with respect to $N$, posing scalability concerns; (2) It is crucial that the text and image in each individual sample remains tightly coupled.
@@
The Context Perceiver take the lengthy outputs from either the Vision Perceiver or the Language Perceiver as input and produces a condensed set of contextual representations. This is achieved by using a fixed number of learnable embeddings of size $\mathbb{R}^{\frac{M}{2}\times d}$ as the input of the Context Perceiver, where $M$ is typically significantly smaller than $L$. These embeddings undergo processing through a self-attention layer, followed by a cross-attention layer with the feature matrix from either the Vision Perceiver or the Language Perceiver, and a feedforward layer. Specifically, for the Vision Perceiver, the learnable embeddings are denoted as $\mathbf{H}^{\text{IMG}}\in\mathbb{R}^{\frac{M}{2}\times d}$, and the output after applying the Context Perceiver is denoted as:
@@ 3.2 Model Training
where $\theta\leftarrow(\theta^{\text{LLM}},\theta^{\text{VP}},\theta^{\text{LP}},\theta^{\text{CP}},\mathbf{H})$ is the model trainable parameter ($\theta^{\text{LLM}}$ is the the parameter of LLM); $y_{i}$ is the ground-truth target, and $\hat{y}_{1:i-1}$ is the $i-1$ preceding tokens of output $y_{i}$; Specifically, we use Vicuna-7B and Vicuna-13B as our backbone language model, resulting in two models CaMML-7B and CaMML-13B.
@@ 4.1 Multimodal Reasoning on ScienceQA
Method AVG. IMG TXT
Human Average 88.40 87.50 89.60
UnifiedQA [36] 74.11 66.53 66.42
GPT-3 CoT [36] 75.17 67.43 74.68
GPT-4 CoT [38] 83.99 71.49 82.65
LLaMA-Adapter [61] 85.19 80.32 83.72
[…]
LLaVA-7B [32] 89.28 87.32 90.96
LLaVA-13B [32] 90.90 88.00 89.49
CaMML-7B 91.32 89.24 93.21
CaMML-13B 92.03 89.94 93.84
[…]
IDEFICS-9B [25] LLaMA-7B 353M+1M 50.9 38.4 35.5 - 25.9 - - 48.2 25.2 - -
InstructBLIP [12] Vicuna-7B 129M+1.2M - 49.2 34.5 60.5 50.1 - - 36.0 23.7 53.4 26.2
Qwen-VL [6] Qwen-7B 1.4B+50M 78.8 59.3 35.2 67.1 63.8 - - 38.2 7.4 56.3 -
Qwen-VL-Chat [6] Qwen-7B 1.4B+50M 78.2 57.5 38.9 68.2 61.5 - 1487.5 60.6 56.7 58.2 -
LLaVA-1.5 [31] Vicuna-7B 558K+665K 78.5 62.0 50.0 66.8 58.2 85.9 1510.7 64.3 58.3 58.6 30.5
CaMML-7B Vicuna-7B ${}^{\dagger}$+665K 79.4 62.7 51.2 67.9 58.0 86.4 1506.9 66.9 60.6 60.4 32.2
IDEFICS-80B [25] LLaMA-65B 353M+1M 60.0 45.2 36.0 - 30.9 - - 54.5 38.1 - -
BLIP-2 [28] Vicuna-13B 129M 41.0 41.0 19.6 61.0 42.5 85.3 1293.8 - - 46.4 22.4
InstructBLIP [12] Vicuna-13B 129M+1.2M - 49.5 33.4 63.1 50.7 78.9 1212.8 - - - 25.6
Shikra [8] Vicuna-13B 600K+5.5M 77.4 - - - - - - 58.8 - - -
LLaVA-1.5 [31] Vicuna-13B 558K+665K 80.0 63.3 53.6 71.6 61.3 85.9 1531.3 67.7 63.6 61.6 35.4
CaMML-13B Vicuna-13B ${}^{\dagger}$+665K 80.2 63.7 57.4 72.3 59.9 86.7 1588.7 70.2 63.6 62.3 36.4
[…]
ScienceQA, as introduced in [36], serves as a multimodal benchmark designed for the task of science question answering. It comprises 21,000 multiple-choice questions, encompassing various domains including biology, physics, chemistry, Earth science, and more. We use the ScienceQA train split to build the datastore search index and for the model training. In specific, Vicuna-v1.3 is used as the LLM backbone, and CLIP-ViT-L-14 [43] is adopted as the visual feature encoder. By default, we retrieve three samples from the datastore as the contexts. We conduct evaluation of the ScienceQA on test split and adopt the same datastore (i.e., train split) search index for contextual samples retrieval.
[…]
@@
 RA-CM3 [56] (2) 89.1 - - - -
 ReViLM [55] (0) 60.8 52.1 - - -
 ReViLM [55] (2) 77.2 - - - -
 ReViLM [55] (4) 90.5 - - - -
 ReViLM [55] (8) 90.2 - - - -
@@
 KOSMOS-1 (4) [17] 101.7 75.3 - 51.8 35.3
 MMICL (4) [64] - 72.0 - 70.6 50.3
@@
Table 3: Comparison with state-of-the-art zero/few-shot models including retrieval-augmented counterparts on Captioning and VQA tasks: CaMML models are the ones that trained in instruction tuning (Sec 4.2) and CaMML are evaluated with 3-shots. CaMML-13B achieves the best performance on 5/5 tasks, even outperforming Flamingo-80B model [3] under 32 shots.
4.2 Multimodal Instruction Tuning
[…]
We compare CaMML with large multimodal baselines such as BLIP-2 [28], InstructBLIP [12], Shikra [8], IDEFICS-9B [25], IDEFICS-80B [25], Qwen-VL [6], Qwen-VL-Chat [6], and LLaVA-1.5 [31]. As shown in Table 2, CaMML-13B demonstrates the best performance across all 11 benchmarks, surpassing LLaVA-1.5, which uses the same training data. Notably, when employing Vicuna-13B as its backbone, CaMML-13B outperforms BLIP-2, InstructBLIP-13B, Shikra, and LLaVA-1.5-13B. Additionally, CaMML-7B delivers impressive results, outperforming all other baseline models except for LLaVA-1.5-13B on ten of the benchmark datasets. On POPE, CaMML-7B achieves performance on par with that of CaMML-13B. These encouraging observations showcase CaMML’s efficacy in harnessing multimodal contextual information and following multimodal instructions. In contrast to Liu et al. [31]’s assertion that LLaVA-1.5 cannot handle multiple images due to the absence of instruction-tuning data and context length limitations, our design demonstrates feasibility of training a model to effectively process multiple images merely with the same training set and efficiently manage long context length.
4.3 Multimodal Few-shot Learning
CaMML essentially operates as a few-shot learning method. Few-shot models leverage the synergy of multiple intertwined images and texts to guarantee precise inference. Among these, retrieval-augmented models are widely recognized for their ability to retrieve highly similar samples, thereby improving few-shot prediction. In Table 3, we conduct a comparison between CaMML, retrieval-augmented models (RA-CM3 [56] and ReViLM [55]), and other popular few-shot LLMs such as Flamingo [3], KOSMOS-1 [17], MMICL [64] on tasks including visual captioning (COCO caption, Flickr30k), and visual question answering (OKVQA, VQAv2, Vizwiz). For CaMML, we employ the same model described in Section 4.2, with LLaVA-665K serving as the datastore.
@@ 4.3 Multimodal Few-shot Learning
 - w/o Perceiver 89.7 85.8 93.2
 - w/o Vision Perceiver 89.8 87.4 92.1
 - w/o Language Perceiver 90.0 87.7 92.0
 - w/o Shared weights 91.3 89.1 93.2
Table 4: Ablation Experiments on CaMML perceiver components and the existence of shared-weights Context Perceiver: CaMML-7B alternatives are evaluated on ScienceQA-test.

### v3

@@ Abstract
In this work, we introduce Context-Aware MultiModal Learner (CaMML), for tuning large multimodal models (LMMs). CaMML, a lightweight module, is crafted to seamlessly integrate multimodal contextual samples into large models, thereby empowering the model to derive knowledge from analogous, domain-specific, up-to-date information and make grounded inferences. Importantly, CaMML is highly scalable and can efficiently handle lengthy multimodal context examples owing to its hierarchical design. Based on CaMML, we have developed two multimodal models, CaMML-7B and CaMML-13B, that have shown exceptional performance across an array of benchmark datasets for multimodal tasks. Remarkably, CaMML-13B achieves the state-of-the-art performance on over ten widely recognized multimodal benchmark datasets, surpassing LLaVA-1.5 (13B) with a noticeable margin, without integration of any external resources. Moreover, we have conducted extensive ablative studies to inspect the inner workings of CaMML and performed qualitative analyses to showcase its effectiveness in handling real-world challenging cases. Code and models are available at: https://github.com/amazon-science/camml.
CaMML: Context-Aware Multimodal Learner for Large Models
Yixin Chen The Chinese University of Hong Kong yxchen@cse.cuhk.edu.hk                        Shuai Zhang Amazon Web Services shuaizs@amazon.com
Boran Han Amazon Web Services boranhan@amazon.com                        Tong He Amazon Web Services htong@amazon.com                        Bo Li The University of Chicago bol@uchicago.edu
@@ 1 Introduction
Recently, large multimodal models (LMMs) Liu et al. (2023b); Zhang et al. (2023a); Wang et al. (2022); Liu et al. (2023a); Zhu et al. (2023); Tsimpoukelli et al. (2021); Alayrac et al. (2022) have demonstrated remarkable performance in a variety of tasks, including but not limited to visual question answering, image captioning, visual grounding, visual-language reasoning, optical character recognition, and visual entailment. Notably, in certain benchmark assessments, these multimodal foundation models have even exceeded human-level performance Zhang et al. (2023b); Liu et al. (2023b).
Despite the impressive performance, their ability to make inferences is constrained by the knowledge encoded in the model parameters. The inflexible design of these models makes it challenging for them to generalize from contextual examples. For instance, LLaVA-1.5 falls short in processing multiple images and attributes it to the lack of corresponding instruction-tuning training data Liu et al. (2023a). Nevertheless, learning and making inferences through contextual examples are fundamental elements of our cognitive processes. Human beings frequently tackle intricate problems by relying on past experiences and identifying analogous situations. Taking inspiration from the cognitive process, we hypothesize that empowering large multimodal models with the capability to perceive and derive insights from analogous contextual examples can significantly streamline the inference process and lead to more precise predictions. Nonetheless, the means to replicate the human cognitive processes in LMMs remain unclear.
As such, our goal is to empower multimodal foundational models to harness context-aware learning, thereby enhancing their ability to comprehend and adapt to previously unseen examples. Identifying relevant context examples is relatively straightforward; this process can be facilitated using multimodal embedding models like ImageBind Girdhar et al. (2023) or CLIP Radford et al. (2021). This approach simulates the act of recalling similar situations from past experiences. Yet, effectively and efficiently integrating the identified similar context samples into large models poses challenges, particularly given the potential variability in the number of context samples and interleaved modalities, resulting in lengthy and heterogeneous context input.
@@ 2 Related Work
The success of LLMs has sparked a burgeoning interest in scaling up multimodal models. One prevalent strategy involves the incorporation of vision encoders, such as ViT Dosovitskiy et al. (2021) and CLIP, into existing LLMs (e.g., LLaMA Touvron et al. (2023), Vicuna Zheng et al. (2023)). For example, LLaMA-adapter Zhang et al. (2023a) inserts learnable adaption prompts into LLaMA and make it to follow multimodal instructions for multimodal reasoning. BLIP2 Li et al. (2023b) bridges different modalities via Q-Former and employs a two-stage training method to bootstrap both representation learning and LLM (e.g., OPT Zhang et al. (2022), Flan-T5 Chung et al. (2022)) generative learning capabilties. LLaVA Liu et al. (2023b, a) projects encoded image features to text token space using linear layers and shows state-of-the-art performance on a variety of multimodal tasks. Another noteworthy series of strategy is to unify multimodal input data and pretrain the model from scratch, exemplified by OFA Wang et al. (2022), Perceiver Jaegle et al. (2021, 2022), Uni-Perceiver Zhu et al. (2022), Unival Shukor et al. (2023), and Unified-IO Lu et al. (2023a). These models map images, text, and other modalities into the same IO space and use a causal language model objective for training.
@@
Extensive research has been conducted on learning from a limited number of multimodal examples. Among them, Flamingo Alayrac et al. (2022) introduces gated xatten-dense layers to establish cross-modal interactions between visual input and text input for few-shot learning. Frozen Tsimpoukelli et al. (2021) trains a vision encoder to produce a sequence of image embeddings and input it to frozen language models for multimodal tasks. An alternative approach for tackling multimodal few-shot learning involves leveraging retrieval augmented generations (RAG) Lewis et al. (2020); Borgeaud et al. (2022); Ram et al. (2023); Khandelwal et al. (2020). This technique enables the models to access external knowledge repositories, databases, or structured data sources, allowing them to access the up-to-date as well as domain-specific expertise when crafting responses. Recently, there has been a surging interest for harnessing RAG within the realm of multimodal models Chen et al. (2022); Yang et al. (2023); Yasunaga et al. (2023); Liu et al. (2023c). Amalgamating RAG with multimodal models helps provide contextually-rich responses and opens up exciting possibilities across domains such as image captioning Yang et al. (2023), image generation Yasunaga et al. (2023), etc. Among them, Chen et al. (2022) propose MuRAG, a model of size 527M, for multimodal-QA. MuRAG concatenates the visual embeddings (extracted by ViT) and text word embeddings of retrieved multimodal image-text pairs and requires the language model to handle a very lengthy input sequences. Re-ViLM Yang et al. (2023) is tailored for image captioning and it follows the Flamingo architecture and only uses the retrieved text for augmentation. RA-CM3 Yasunaga et al. (2023) leverages retrieval to augment the CM3 Aghajanyan et al. (2022) pipeline which can do both text-to-image and image-to-text generations.
In contrast to previous approaches, our primary objective is to develop a lightweight module capable of processing multimodal context information efficiently and effectively, with superior generalization capabilities. This novel approach enables LLM to proficiently reason from multimodal in-context examples, leading to more accurate and precise inferences. As discussed in Liu et al. (2023a), LLaVA-1.5 faces limitations when confronted with scenarios involving multiple images and lengthy contexts. CaMML can effectively address these challenges, a fact further corroborated by its superior performances across multiple benchmarks.
@@ 3.1.1 Datastore and Context Retriever
The datastore is created from either the training set or external resources. We adopt an embedding encoder, ImageBind Girdhar et al. (2023), to extract dense vector representations for every multimodal sample contained in the datastore. Following this, we build an index using Faiss Johnson et al. (2019a), a highly efficient similarity search library, to enable rapid and efficient search operations. During both the training and inference phases, our approach involves identifying the top $N$ most closely related multimodal samples, denoted as $C_{1},C_{2},\ldots,C_{N}$, from the datastore using Faiss, based on a given multimodal query $q$ [^1] Specifically, we compute the similarity between query text and context image, and top-k images along with their corresponding texts are recalled from the datastore.. It is worth noting that the retriever remains frozen throughout all stages of the process.
@@ 3.1.2 Multimodal CaMML Perceiver
To seamlessly integrate interleaved multimodal samples, we introduce a novel module denoted as CaMML Perceiver, which revolves around two key design principles: (1) The module should accommodate a dynamic number of context samples without imposing a significant computational burden. It is worth noting that directly concatenating the image tokens and text tokens will result in $L=\sum^{N}_{i=1}(T_{i}^{\text{IMG}}+T_{i}^{\text{TXT}})$ tokens, where $T_{i}^{\text{IMG}}$ is the number of image tokens and $T_{i}^{\text{TXT}}$ is the number of text tokens. The formulation results in a linear increase of $L$ with respect to $N$, posing scalability concerns; (2) It is crucial that the text and image in each individual sample remain tightly coupled.
@@
The Context Perceiver takes the lengthy outputs from either the Vision Perceiver or the Language Perceiver as input and produces a condensed set of contextual representations. This is achieved by using a fixed number of learnable embeddings of size $\mathbb{R}^{\frac{M}{2}\times d}$ as the input of the Context Perceiver, where $M$ is typically significantly smaller than $L$. These embeddings undergo processing through a self-attention layer, followed by a cross-attention layer with the feature matrix from either the Vision Perceiver or the Language Perceiver, and a feedforward layer. Specifically, for the Vision Perceiver, the learnable embeddings are denoted as $\mathbf{H}^{\text{IMG}}\in\mathbb{R}^{\frac{M}{2}\times d}$, and the output after applying the Context Perceiver is denoted as:
@@
Method AVG. IMG TXT
Human Average 88.40 87.50 89.60
UnifiedQA Lu et al. (2022a) 74.11 66.53 66.42
GPT-3 CoT Lu et al. (2022a) 75.17 67.43 74.68
GPT-4 CoT Lu et al. (2023b) 83.99 71.49 82.65
LLaMA-Adapter Zhang et al. (2023a) 85.19 80.32 83.72
MMCoTBase Zhang et al. (2023b) 84.91 82.90 87.88
MMCoTLarge Zhang et al. (2023b) 91.68 88.80 95.26
LLaVA-7B Liu et al. (2023b) 89.28 87.32 90.96
LLaVA-13B Liu et al. (2023b) 90.90 88.00 89.49
CaMML-7B 91.32 89.24 93.21
CaMML-13B 92.03 89.94 93.84
@@ 3.2 Model Training
where $\theta\leftarrow(\theta^{\text{LLM}},\theta^{\text{VP}},\theta^{\text{LP}},\theta^{\text{CP}},\mathbf{H})$ is the model trainable parameter ($\theta^{\text{LLM}}$ is the parameter of LLM); $y_{i}$ is the ground-truth target, and $\hat{y}_{1:i-1}$ is the $i-1$ preceding tokens of output $y_{i}$; Specifically, we use Vicuna-7B and Vicuna-13B as our backbone language model, resulting in two models CaMML-7B and CaMML-13B.
@@ 4.1 Multimodal Reasoning on ScienceQA
Method LLM Data VQAv2 GQA VizWiz SQAI VQAT POPE MME MMB MMBCN SEED MM-Vet
IDEFICS-9B Laurençon et al. (2023) LLaMA-7B 353M+1M 50.9 38.4 35.5 - 25.9 - - 48.2 25.2 - -
InstructBLIP Dai et al. (2023) Vicuna-7B 129M+1.2M - 49.2 34.5 60.5 50.1 - - 36.0 23.7 53.4 26.2
Qwen-VL Bai et al. (2023) Qwen-7B 1.4B+50M 78.8 59.3 35.2 67.1 63.8 - - 38.2 7.4 56.3 -
Qwen-VL-Chat Bai et al. (2023) Qwen-7B 1.4B+50M 78.2 57.5 38.9 68.2 61.5 - 1487.5 60.6 56.7 58.2 -
LLaVA-1.5 Liu et al. (2023a) Vicuna-7B 558K+665K 78.5 62.0 50.0 66.8 58.2 85.9 1510.7 64.3 58.3 58.6 30.5
CaMML-7B Vicuna-7B 558K+665K 79.4 62.7 51.2 67.9 58.0 86.4 1506.9 66.9 60.6 60.4 32.2
IDEFICS-80B Laurençon et al. (2023) LLaMA-65B 353M+1M 60.0 45.2 36.0 - 30.9 - - 54.5 38.1 - -
BLIP-2 Li et al. (2023b) Vicuna-13B 129M 41.0 41.0 19.6 61.0 42.5 85.3 1293.8 - - 46.4 22.4
InstructBLIP Dai et al. (2023) Vicuna-13B 129M+1.2M - 49.5 33.4 63.1 50.7 78.9 1212.8 - - - 25.6
Shikra Chen et al. (2023) Vicuna-13B 600K+5.5M 77.4 - - - - - - 58.8 - - -
LLaVA-1.5 Liu et al. (2023a) Vicuna-13B 558K+665K 80.0 63.3 53.6 71.6 61.3 85.9 1531.3 67.7 63.6 61.6 35.4
CaMML-13B Vicuna-13B 558K+665K 80.2 63.7 57.4 72.3 59.9 86.7 1588.7 70.2 63.6 62.3 36.4
Table 2: Comparison with state-of-the-art large multimodal models on 11 benchmarks (VQAv2 Antol et al. (2015), GQA Hudson and Manning (2019), VizWiz Gurari et al. (2018), SQAI Lu et al. (2022a): ScienceQA-IMG, VQAT Singh et al. (2019): TextVQA, POPE Li et al. (2023c), MME Fu et al. (2023), MMB Liu et al. (2023d): MMBench, MMBCN Liu et al. (2023d): MMbench-Chinese, SEED Li et al. (2023a): SEED-Bench, MM-Vet Yu et al. (2023)). CaMML achieves the best performance on 10/11 tasks. denotes BLIP558K-pretrained projector is initialized for instruction tuning.
ScienceQA, as introduced in Lu et al. (2022a), serves as a multimodal benchmark designed for the task of science question answering. It comprises 21,000 multiple-choice questions, encompassing various domains including biology, physics, chemistry, Earth science, and more. We use the ScienceQA train split to build the datastore search index and for the model training. In specific, Vicuna-v1.3 is used as the LLM backbone, and CLIP-ViT-L-14 Radford et al. (2021) is adopted as the visual feature encoder. By default, we retrieve three samples from the datastore as the contexts. We conduct evaluation of the ScienceQA on test split and adopt the same datastore (i.e., train split) search index for contextual samples retrieval.
We compare CaMML with various baselines including UnifiedQA Lu et al. (2022b), GPT-4 CoT Lu et al. (2023b), LLaMA-Adapter Zhang et al. (2023a), MMCoTBase Zhang et al. (2023b), MMCoTLarge Zhang et al. (2023b), LLaVA Liu et al. (2023b) 7B and 13B. As shown in Table 1, CaMML surpass baselines by a noticeable margin. It is worth noting that our model attains the best performance on average, surpassing the previous state-of-the-art model MMCoTLarge [^2] CaMML-13B attains the highest AVG scores on the ScienceQA leaderboard when GPT4-review is not used. We also notice that, on the IMG questions, CaMML-13B outperforms LLaVA-13B and MMCoTLarge by a wide margin, which underscores CaMML’s strong capability in handling challenging questions that incorporate scientific images. Furthermore, we conduct detailed ablative studies with the ScienceQA dataset using CaMML. See Section 5.1 for more details.
@@
RA-CM3 Yasunaga et al. (2023) (2) 89.1 - - - -
ReViLM Yang et al. (2023) (0) 60.8 52.1 - - -
ReViLM Yang et al. (2023) (2) 77.2 - - - -
ReViLM Yang et al. (2023) (4) 90.5 - - - -
ReViLM Yang et al. (2023) (8) 90.2 - - - -
@@
KOSMOS-1 (4) Huang et al. (2023) 101.7 75.3 - 51.8 35.3
MMICL (4) Zhao et al. (2023) - 72.0 - 70.6 50.3
@@
Table 3: Comparison with state-of-the-art zero/few-shot models including retrieval-augmented counterparts on Captioning and VQA tasks: CaMML are the ones that trained in instruction tuning (Sec 4.2) and are evaluated with 3-shots. CaMML-13B achieves the best performance on 5/5 tasks, even outperforming Flamingo-80B model Alayrac et al. (2022) under 32 shots.
@@
- w/o Perceiver 89.7 85.8 93.2
- w/o Vision Perceiver 89.8 87.4 92.1
- w/o Language Perceiver 90.0 87.7 92.0
- w/o Shared weights 91.3 89.1 93.2
Table 4: Ablation Experiments on CaMML perceiver components and shared-weights Context Perceiver: CaMML-7B models are evaluated on ScienceQA-test.
4.2 Multimodal Instruction Tuning
CaMML can be enhanced by tuning to follow multimodal instructions to handle a more diverse set of multimodal tasks. To substantiate this, we perform instruction tuning following Liu et al. (2023a) on LLaVA-665K, an instruction-tuning data mixture of LLaVA-1.5. Following Liu et al. (2023a), we use Vicuna-v1.5 as the LLM backbone and CLIP-ViT-L-14-336px as the visual encoder. The datastore is built using the exact training set, LLaVA-665K, without external injected information for a fair comparison. LLaVA-665K comprises 665K multimodal samples, and is made up of LLaVA-Instruct-158K, ShareGPT-40K sha (2023), VQAv2 Antol et al. (2015), GQA Hudson and Manning (2019), OKVQA Marino et al. (2019), OCRVQA Mishra et al. (2019), A-OKVQA Schwenk et al. (2022), TextCaps Sidorov et al. (2020), RefCOCO Yu et al. (2016) and VG Krishna et al. (2017). We adopt a training strategy in which each individual sample might have a distinct $N$, spanning a range from $1$ to $3$ and denote it as “mixed-shots training". During inference, we retrieve three relevant supporting multimodal samples from the datastore index. To evaluate the model effectiveness, we traverse 11 comprehensive benchmarks, including VQAv2, GQA, TextVQA Singh et al. (2019), MME Fu et al. (2023), POPE Li et al. (2023c), MM-Vet Yu et al. (2023), MMBench Liu et al. (2023d), MMBench-CN Liu et al. (2023d), SEED-Bench Li et al. (2023a), and Vizwiz Gurari et al. (2018). Results on ScienceQA image (denoted as SQAI) is also reported following Liu et al. (2023a).
We compare CaMML with large multimodal baselines such as BLIP-2 Li et al. (2023b), InstructBLIP Dai et al. (2023), Shikra Chen et al. (2023), IDEFICS-9B Laurençon et al. (2023), IDEFICS-80B Laurençon et al. (2023), Qwen-VL Bai et al. (2023), Qwen-VL-Chat Bai et al. (2023), and LLaVA-1.5 Liu et al. (2023a). As shown in Table 2, CaMML-13B demonstrates the best performance across all 11 benchmarks, surpassing LLaVA-1.5, which uses the same training data. Notably, when employing Vicuna-13B as its backbone, CaMML-13B outperforms BLIP-2, InstructBLIP-13B, Shikra, and LLaVA-1.5-13B. Additionally, CaMML-7B delivers impressive results, outperforming all other baseline models except for LLaVA-1.5-13B on ten of the benchmark datasets. On POPE, CaMML-7B achieves performance on par with that of CaMML-13B. These encouraging observations showcase CaMML’s efficacy in harnessing multimodal contextual information and following multimodal instructions. In contrast to Liu et al. (2023a)’s assertion that LLaVA-1.5 cannot handle multiple images due to the absence of instruction-tuning data and context length limitations, our design demonstrates feasibility of training a model to effectively process multiple images merely with the same training set and efficiently manage long context length.
4.3 Multimodal Few-shot Learning
CaMML essentially operates as a few-shot learning method. Few-shot models leverage the synergy of multiple intertwined images and texts to guarantee precise inference. Among these, retrieval-augmented models are widely recognized for their ability to retrieve highly similar samples, thereby improving few-shot prediction. In Table 3, we conduct a comparison between CaMML, retrieval-augmented models (RA-CM3 Yasunaga et al. (2023) and ReViLM Yang et al. (2023)), and other popular few-shot LLMs such as Flamingo Alayrac et al. (2022), KOSMOS-1 Huang et al. (2023), MMICL Zhao et al. (2023) on tasks including visual captioning (COCO caption, Flickr30k), and visual question answering (OKVQA, VQAv2, Vizwiz). For CaMML, we employ the same model described in Section 4.2, with LLaVA-665K serving as the datastore.